Yet if I do intuitive F2L, does that make cuBerBruce's numbers more accurate as probabilities?
To my understanding, intuitive F2L is a general approach to solving F2L, so I think it's hard to draw a conclusion as to how it would affect the actual probabilities. Intuitive F2L doesn't imply any particular manner of selecting which pair to solve next, for instance.
I also note that if you have two pairs left, and then solve one more pair, you would generally expect to have one slot left. But (assuming the 150 possible configurations for the last corner and edge are equally likely), you would have the following probabilities:
Both in top layer: (4/5)*(4/5) = .64
Corner in top layer, edge in slot: (4/5)*(1/5) = .16
Corner in slot, edge in top layer: (1/5)*(4/5) = .16
Both in slot, but at least one mis-oriented: (6-1)/150 = .03333...
Both in slot and correctly oriented (i.e, solved): 1/150 = .00666...
So the probability that the last slot is not solved and at least one of the pieces is not in the top layer is .16 + .16 + .03333... = .35333...
The probability that at least one of the pieces for the last slot is not in the top layer is .36, as badmephisto had calculated. But note that this includes the case of the last slot being solved. Of course, if the last slot is solved, you don't have to move the pieces to the top layer, then solve them again.
I note that in my data, the 212/596 (or 53/149) is the probability for having one unsolved slot (and you know that this last slot is not inadvertantly solved) and at least one piece not in the top layer. The 212/597 figure assumes any of the four slots could be left (except the case where no slot is left, of course), but since you know what slots you solve along the way, I think it's really incorrect to count the configurations for all four cases as far as which slot is the last one. So the 212/597 should really be 54/150 (36%), or 53/150 if you want to count the last slot solved as a possible configuration, but not count it as a "bad" case.
Basically, if you know that n slots are already solved, then configurations where any of those n slots are unsolved should be counted as 0% likely, rather than equally likely. So I would consider the figures my table for 3 or less, 2 or less, 1 or less to be wrong, and I think exactly 3 left and exactly 2 left might also be wrong.